# Low-Resource Domain Adaptation for Speech LLMs via Text-Only Fine-Tuning

### v1

@@ Abstract
Index Terms:
speech large language model, automatic speech recognition, unpaired text fine-tune, domain adaptation
@@ I Introduction
Automatic speech recognition (ASR) is a core technology in intelligent speech systems. It has evolved from traditional ASR models [1, 2, 3] to end-to-end (E2E) models [4, 5, 6, 7, 8] and is now undergoing a third paradigm shift toward Speech LLMs [9, 10]. Speech LLMs extend general-purpose LLMs by incorporating audio encoders that extract acoustic features, which are then projected into the LLM input space and combined with text token embeddings [11, 12]. Empowered by the strong contextual reasoning capabilities of LLMs, Speech LLMs have demonstrated remarkable performance in the field of ASR. Recent models such as SeedASR [13] and FireRedASR [14] have achieved state-of-the-art results on benchmark datasets including LibriSpeech [15] and AISHELL [16] , significantly surpassing the performance of previous approaches.
Leveraging the strong learning capacity of Speech LLMs, these models have achieved outstanding recognition performance on widely-used benchmark datasets. However, this success also introduces a significant risk: the training and performance improvements of Speech LLMs heavily depend on the availability of large-scale data. In scenarios where training data are sparse, the model is prone to overfitting to specific domains[17], resulting in degraded generalization to unseen or broader domains. And a more important question is, in low-resource speech domains, the scarcity of paired audio-text data[18, 19] poses a major bottleneck, making it difficult to perform domain-specific fine-tuning and leading to poor recognition accuracy in those domains.
One approach leverages the relative abundance of textual data in low-resource domains. Given sufficient in-domain text, an LLM can generate domain-specific content, and a text-to-speech (TTS) system can synthesize corresponding audio to create pseudo audio-text pairs. However, synthetic speech often lacks naturalness and variability [20, 21], and the overall pipeline incurs significant computational and engineering costs. Another direction reduces the number of trainable parameters in Speech LLMs[22, 23], such as prompt tuning or prefix tuning. While these methods lessen retraining overhead, they require domain-specific prompt design, limiting generalizability and introducing additional manual effort.
@@ II-B Domain Adaptation For End to End Models
In end-to-end (E2E) models, domain adaptation can also be performed by replacing the LM; however, the LM is no longer a necessary component of E2E architectures. In such systems, the LM is typically incorporated via shallow fusion [24], deep fusion [25], or cold fusion [26]. For deep fusion and cold fusion approaches, domain adaptation through LM replacement is challenging because the LM is implicitly integrated into the model. The recognition probability in this case can be formulated as
@@
Another approach is to augment the model with relevant information from the target domain [27]. For example, Pundak et al. [28] incorporate contextual information through a context encoder and an attention mechanism. Huang et al. [29] propose combining context embeddings with acoustic representations extracted from the acoustic encoder and text representations obtained using a prediction network based on previous sub-word units. Yang et al. [30] further propose an effective method to construct high-quality context lists for unified streaming and non-streaming E2E models. However, these methods require additional resources or supervision.
@@ II-C Domain Adaptation For Synthetic Speech
The development of synthetic speech technology enables the use of speech synthesized from text in data-scarce domains to promote domain adaptation. Su et al. [31] studied an E2E pipeline where LLMs are used to generate text, followed by speech synthesis and domain adaptation, potentially eliminating the reliance on textual data. D et al. [32] adopt the method proposed by Thomas et al. [25] to investigate whether synthetic speech can be substituted by injected contextual information.
@@ II-D Domain Adaptation For Speech LLMs
Speech large language models (Speech LLMs) leverage extensive general-domain textual information acquired during pre-training and can further inject domain-specific knowledge by exploiting their contextual capabilities. Yang et al. [33] improve recognition performance by prompting the LLM with hot words extracted from slide text. Similarly, Lakomkin et al. [34] incorporate metadata such as titles to enhance recognition accuracy. Nevertheless, obtaining specific contextual information for each speech utterance remains challenging.
Li et al. [22] propose using domain-specific text prompts to improve recognition performance in a particular domain; however, the prompts must be carefully designed, and devising a stable and broadly applicable prompting strategy remains an open problem. Liao et al. [23] investigate using LLMs to generate descriptions for corresponding speech segments, allowing models like Whisper to leverage these descriptions for recognition enhancement. Despite its potential, generating a text description for each utterance is often impractical.
@@ III-A Fine-Tuning LLM on Target Domain Text for Adaptation
where $\mathrm{LLM}_{\theta_{s}+\phi}$ denotes a large language model initially pre-trained on general-domain text $\theta_{s}$ and subsequently adapted to the source domain via cross-modal training with paired speech-text data. The model parameters, including the pretrained encoder $\mathrm{Encoder}_{\phi}$, the projection module $\mathrm{Projector}_{\phi}$, and the source-domain adapted LLM $\mathrm{LLM}_{\theta_{s}+\phi}$ , are jointly optimized by minimizing the cross-entropy loss between the predicted transcription and the ground-truth text. We refer to this stage as source domain pretraining, as illustrated on the left side of Fig. 1. Both the projector and the LLM are included in the training process to achieve better generalization ability [17].
@@ IV-A Dataset
Table I summarizes the datasets used. LibriSpeech [15] (1,000 hours) serves as the source domain, comprising audiobook recordings with test-clean and test-other subsets. For cross-domain evaluation, we adopt SlideSpeech [35] (470 hours) from the online meeting domain and Medical [36] (8 hours) from the healthcare domain, simulating diverse low-resource scenarios across academic and specialized medical fields. Additionally, GigaSpeech [37] (10,000+ hours) provides large-scale evaluation for generalization assessment. These datasets collectively support rigorous cross-domain adaptation experiments.
@@ IV-B Model Architecture
As illustrated in Figure 1, our Speech LLM consists of a speech encoder, a projector, and an LLM decoder. We adopt Whisper-large-v3 [38] as the encoder, with 635M parameters and 128-dimensional MFCC features, padding all inputs to 30 seconds. The encoder has a downsampling factor of 2, yielding a fixed output of 1500 frames with 1280 dimensions.
@@
For the decoder, we use Qwen2.5-7B-Instruct [39], a 7B-parameter LLM. We fine-tune it using LoRA, applying modifications to the query, key, value, and output projection matrices in each decoder layer, while freezing the rest. LoRA is configured with a rank of 64, dropout of 5%, and $\text{LoRA}_{\alpha}$ of 16, introducing 161M trainable parameters.
@@ IV-C Training Strategy
In the source domain pretraining stage, the training proceeds in two phases: first optimizing the projector for 3–5 epochs until convergence, followed by joint optimization of the projector and LoRA parameters for 1–2 epochs, following the main training strategy such as SLAMOON [40] and OSUM [12].
@@ V-A Generalization Evaluation of Speech LLM After Source Domain Pretraining
As shown in the first row of Table III, the base model is trained on LibriSpeech as the source domain. Consistent with findings in [17], the Speech LLM achieves relatively good performance on LibriSpeech due to its large parameter size and superior learning capacity. (Note that with more training epochs, the performance could be further improved on source dataset; however, we limit the epochs for efficient experimentation and relatively good generalization ability.) However, its performance significantly degrades on cross-domain datasets (SlideSpeech, GigaSpeech, etc.), revealing critical limitations in domain generalization.

### v2

@@ I Introduction
Automatic speech recognition (ASR) is a core technology in intelligent speech systems. It has evolved from traditional ASR models [levinson1983introduction, lee1989speaker, 9536732] to end-to-end (E2E) models [abdel2014convolutional, gulati2020conformer, gao2022paraformer, graves2005framewise, ren2022improving] and is now undergoing a third paradigm shift toward Speech LLMs [peng2025surveyspeechlargelanguage, arora2025landscapespokenlanguagemodels]. Speech LLMs extend general-purpose LLMs by incorporating audio encoders that extract acoustic features, which are then projected into the LLM input space and combined with text token embeddings [ma2024embarrassingly, tang2023salmonn]. Empowered by the strong contextual reasoning capabilities of LLMs, Speech LLMs have demonstrated remarkable performance in the field of ASR. Recent models such as SeedASR [bai2024seed] and FireRedASR [xu2025fireredasr] have achieved state-of-the-art results on benchmark datasets including LibriSpeech [panayotovLibrispeechASRCorpus2015] and AISHELL [buAISHELL1OpensourceMandarin2017] , significantly surpassing the performance of previous approaches.
Leveraging the strong learning capacity of Speech LLMs, these models have achieved outstanding recognition performance on widely-used benchmark datasets. However, this success also introduces a significant risk: the training and performance improvements of Speech LLMs heavily depend on the availability of large-scale data. In scenarios where training data are sparse, the model is prone to overfitting to specific domains[kumar2024performance], resulting in degraded generalization to unseen or broader domains. And a more important question is, in low-resource speech domains, the scarcity of paired audio-text data[singhal2023domain, ashok2024domain] poses a major bottleneck, making it difficult to perform domain-specific fine-tuning and leading to poor recognition accuracy in those domains.
One approach leverages the relative abundance of textual data in low-resource domains. Given sufficient in-domain text, an LLM can generate domain-specific content, and a text-to-speech (TTS) system can synthesize corresponding audio to create pseudo audio-text pairs. However, synthetic speech often lacks naturalness and variability [dunbar2019zero, 10389722], and the overall pipeline incurs significant computational and engineering costs. Another direction reduces the number of trainable parameters in Speech LLMs[li2023prompting, liao2023zero], such as prompt tuning or prefix tuning. While these methods lessen retraining overhead, they require domain-specific prompt design, limiting generalizability and introducing additional manual effort.
@@ II-B Domain Adaptation For End to End Models
In end-to-end (E2E) models, domain adaptation can also be performed by replacing the LM; however, the LM is no longer a necessary component of E2E architectures. In such systems, the LM is typically incorporated via shallow fusion [zhao2019shallow], deep fusion [thomas2022integrating], or cold fusion [sriram2017cold]. For deep fusion and cold fusion approaches, domain adaptation through LM replacement is challenging because the LM is implicitly integrated into the model. The recognition probability in this case can be formulated as
@@
Another approach is to augment the model with relevant information from the target domain [aleksic2015bringing]. For example, Pundak et al. [pundak2018deep] incorporate contextual information through a context encoder and an attention mechanism. Huang et al. [huang2023contextualized] propose combining context embeddings with acoustic representations extracted from the acoustic encoder and text representations obtained using a prediction network based on previous sub-word units. Yang et al. [yang2023two] further propose an effective method to construct high-quality context lists for unified streaming and non-streaming E2E models. However, these methods require additional resources or supervision.
@@ II-C Domain Adaptation For Synthetic Speech
The development of synthetic speech technology enables the use of speech synthesized from text in data-scarce domains to promote domain adaptation. Su et al. [su2024corpus] studied an E2E pipeline where LLMs are used to generate text, followed by speech synthesis and domain adaptation, potentially eliminating the reliance on textual data. D et al. [d2023can] adopt the method proposed by Thomas et al. [thomas2022integrating] to investigate whether synthetic speech can be substituted by injected contextual information.
@@ II-D Domain Adaptation For Speech LLMs
Speech large language models (Speech LLMs) leverage extensive general-domain textual information acquired during pre-training and can further inject domain-specific knowledge by exploiting their contextual capabilities. Yang et al. [yang2024mala] improve recognition performance by prompting the LLM with hot words extracted from slide text. Similarly, Lakomkin et al. [lakomkin2024end] incorporate metadata such as titles to enhance recognition accuracy. Nevertheless, obtaining specific contextual information for each speech utterance remains challenging.
Li et al. [li2023prompting] propose using domain-specific text prompts to improve recognition performance in a particular domain; however, the prompts must be carefully designed, and devising a stable and broadly applicable prompting strategy remains an open problem. Liao et al. [liao2023zero] investigate using LLMs to generate descriptions for corresponding speech segments, allowing models like Whisper to leverage these descriptions for recognition enhancement. Despite its potential, generating a text description for each utterance is often impractical.
@@ III-A Fine-Tuning LLM on Target Domain Text for Adaptation
where $\mathrm{LLM}_{\theta_{s}+\phi}$ denotes a large language model initially pre-trained on general-domain text $\theta_{s}$ and subsequently adapted to the source domain via cross-modal training with paired speech-text data. The model parameters, including the pretrained encoder $\mathrm{Encoder}_{\phi}$, the projection module $\mathrm{Projector}_{\phi}$, and the source-domain adapted LLM $\mathrm{LLM}_{\theta_{s}+\phi}$ , are jointly optimized by minimizing the cross-entropy loss between the predicted transcription and the ground-truth text. We refer to this stage as source domain pretraining, as illustrated on the left side of Fig. 1. Both the projector and the LLM are included in the training process to achieve better generalization ability [kumar2024performance].
@@ IV-A Dataset
Table I summarizes the datasets used. LibriSpeech [panayotovLibrispeechASRCorpus2015] (1,000 hours) serves as the source domain, comprising audiobook recordings with test-clean and test-other subsets. For cross-domain evaluation, we adopt SlideSpeech [wangSlideSpeechLargeScale2024] (470 hours) from the online meeting domain and Medical [medical] (8 hours) from the healthcare domain, simulating diverse low-resource scenarios across academic and specialized medical fields. Additionally, GigaSpeech [chenGigaSpeechEvolvingMultidomain2021] (10,000+ hours) provides large-scale evaluation for generalization assessment. These datasets collectively support rigorous cross-domain adaptation experiments.
@@ IV-B Model Architecture
As illustrated in Figure 1, our Speech LLM consists of a speech encoder, a projector, and an LLM decoder. We adopt Whisper-large-v3 [radfordRobustSpeechRecognition2023] as the encoder, with 635M parameters and 128-dimensional MFCC features, padding all inputs to 30 seconds. The encoder has a downsampling factor of 2, yielding a fixed output of 1500 frames with 1280 dimensions.
@@
For the decoder, we use Qwen2.5-7B-Instruct [qwen2.5], a 7B-parameter LLM. We fine-tune it using LoRA, applying modifications to the query, key, value, and output projection matrices in each decoder layer, while freezing the rest. LoRA is configured with a rank of 64, dropout of 5%, and $\text{LoRA}_{\alpha}$ of 16, introducing 161M trainable parameters.
@@ IV-C Training Strategy
In the source domain pretraining stage, the training proceeds in two phases: first optimizing the projector for 3–5 epochs until convergence, followed by joint optimization of the projector and LoRA parameters for 1–2 epochs, following the main training strategy such as SLAMOON [geng2025osum] and OSUM [tang2023salmonn].
@@ V-A Generalization Evaluation of Speech LLM After Source Domain Pretraining
As shown in the first row of Table III, the base model is trained on LibriSpeech as the source domain. Consistent with findings in [kumar2024performance], the Speech LLM achieves relatively good performance on LibriSpeech due to its large parameter size and superior learning capacity. (Note that with more training epochs, the performance could be further improved on source dataset; however, we limit the epochs for efficient experimentation and relatively good generalization ability.) However, its performance significantly degrades on cross-domain datasets (SlideSpeech, GigaSpeech, etc.), revealing critical limitations in domain generalization.
@@ VI Conclusion and Future Work
VII Acknowledgment
This research was carried out by Yangui Fang and Jing Peng while they were interning at AISPEECH. We would like to express our gratitude to AISPEECH for providing the necessary computing resources to support this work.